# On the growth of actions of free products

### v1

@@ 2.3. A subclass closed under graph products
In this section we prove that, under a mild assumption on the function $f$, a certain subclass of $\mathcal{C}(f)$ is closed under finite graph products (Proposition 2.13). The definition of this subclass is inspired by the work of the third author [Sal21], and Proposition 2.13 elaborates on [Sal21, Theorem 3]. In particular Proposition 2.13 provides an elementary proof of the result from [Sal21] that finitely generated right-angled Artin groups admit a faithful action with linear growth. (Note that Theorem 3 in [Sal21] is stronger than that).
@@
Let $C_{1}$ coming from the fact that $f$ satisfies (1). Let $n\geq 1$. Let $x$ be a vertex in $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$, and let $\ell$ such that $x$ belongs to $X_{\ell}$. Let $B(n)$ denote the ball of radius $n$ around $x$ in $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$. Let $X_{\ell-\ell_{1}},\ldots X_{\ell},\ldots X_{\ell+\ell_{2}}$ be the members of $(X_{k})_{k\leq q}$ that $B(n)$ intersects. Certainly we have
@@
Let $G=\langle S\rangle$ be a finitely generated group, and $X$ a $G$-set. We say that the $G$-action on $X$ has large displacement if all $G$-orbits in $X$ are finite, and if there is $C>0$ such that for every non-trivial element $g\in G$, there is a $G$-orbit $O$ in $X$ and $x\in O$ such that $gx\neq x$ and $\mathrm{diam}(\Gamma(G,O))\leq Cd(x,gx)$.
Here by $\Gamma(G,O)$ we mean the Schreier graph of the $G$-action on $O$ (which is a connected component in the larger graph $\Gamma(G,X)$). It is easily checked that this condition does not depend on the choice of $S$ (although $C$ can depend on $S$). Note that the definition forces in particular the group $G$ to be residually finite.
@@
We denote by $\mathcal{C}_{LD}(f)\subset\mathcal{C}(f)$ the class of groups admitting an action with large displacement and of growth at most $f(n)$.
@@
    Every faithful action of a finite group has large displacement.
@@
    The action of $\mathbb{Z}$ on $\bigsqcup_{n\geq 1}\mathbb{Z}/n\mathbb{Z}$ has large displacement. Hence $\mathbb{Z}$ belongs to $\mathcal{C}_{LD}(\mathrm{lin})$.
@@
    If $G_{1},\ldots,G_{k}$ admit actions with large displacement respectively on $X_{1},\ldots,X_{k}$, then the action of $G_{1}\times\cdots\times G_{k}$ on $X_{1}\sqcup\cdots\sqcup X_{k}$ has large displacement. Hence each class $\mathcal{C}_{LD}(f)$ is closed under direct products.
@@
For every $d\geq 1$ and $p\geq 2$, the wreath product $G=(\mathbb{Z}/p\mathbb{Z})\wr\mathbb{Z}^{d}$ belongs to $\mathcal{C}_{LD}(f)$ for $f(n)=n^{d}$.
@@
We denote elements of $G$ as pairs $(f,u)$, with $f\colon\mathbb{Z}^{d}\to\mathbb{Z}/p\mathbb{Z}$ of finite support and $u\in\mathbb{Z}^{d}$. For $m\geq 2$, set $H_{m}=\{(f,u)\colon\quad\sum_{v\in(m\mathbb{Z})^{d}}f(v)=0,u\in m\mathbb{Z}^{d}\}$. Then $H_{m}$ is a finite index subgroup of $G$, of index $pm$. Set $X_{m}:=G/H_{m}$ and $X=\sqcup_{m}X_{m}$. We check that the action of $G$ on $X$ has large displacement and satisfies $\operatorname{vol}_{G,X}(n)\simeq n^{d}$. Consider the standard generating set $S=\{(\delta_{0},0),(0,e_{i}),i=1,\dots d\}$, where $\delta_{0}\colon\mathbb{Z}^{d}\to\mathbb{Z}/p\mathbb{Z}$ takes the value 1 on 00 and 0 elsewhere, and $e_{1},\ldots,e_{d}$ is the standard basis of $\mathbb{Z}^{d}$. Note first that the map $G\to(\mathbb{Z}/p\mathbb{Z})\times(\mathbb{Z}/m\mathbb{Z})^{d}$ that sends $(f,u)\in G$ to $(\sum_{v\in m\mathbb{Z}^{d}}f(v),u\mod m)$ descends to a bijection of the coset space $G/H_{m}$ with $(\mathbb{Z}/p\mathbb{Z})\times(\mathbb{Z}/m\mathbb{Z})^{d}$, so that we may identify $X_{m}$ with $(\mathbb{Z}/p\mathbb{Z})\times(\mathbb{Z}/m\mathbb{Z})^{d}$. Under this identification, the action of $G$ on $X_{m}$ coincides with the standard wreath product action of the natural quotient $(\mathbb{Z}/p\mathbb{Z})\wr(\mathbb{Z}/m\mathbb{Z})^{d}$. Explicitly, the action of elements in the generating set $S$ is described as follows: the lamp generator $(\delta_{0},0)$ permutes cyclically $(\mathbb{Z}/p\mathbb{Z})\times\{0\}$ and acts trivially elsewhere; each element $(0,e_{i})$ maps $(r,u\text{ mod }m)$ to $(r,u+e_{i}\text{ mod }m)$. In particular the Schreier graph of the $G$-action on $X_{m}$ is isomorphic (ignoring loops) to the graph obtained taking a cycle of length $p$ and gluing to each point a copy of the standard Cayley graph of $(\mathbb{Z}/m\mathbb{Z})^{d}$. It follows that $\operatorname{vol}_{G,X}(n)\simeq n^{d}$, and that $C_{1}m\leq\mathrm{diam}(X_{m})\leq C_{2}m$ for some constants $C_{1},C_{2}$. For the large displacement property, let $g=(f,v)$ be a non-trivial element of $G$. Suppose first that $v\neq 0$, and choose $m=2|v|$, where ||$\cdot$|| ⋅ | denotes the standard word metric of $\mathbb{Z}^{d}$. Then, since the natural projection of $X_{m}$ to $(\mathbb{Z}/m\mathbb{Z})^{d}$ is equivariant, for every $x\in X_{m}$ we have $d(x,gx)\geq|v|\geq\frac{1}{2}C_{1}\mathrm{diam}(X_{m})$. Suppose now that $v=0$. Let $u$ be such that $f(u)\neq 0$ and with $|u|$ maximal, and choose again $m>2|u|$. Let $x\in X_{m}\simeq(\mathbb{Z}/p\mathbb{Z})\times(\mathbb{Z}/m\mathbb{Z})^{d}$ be the point $x=(0,u\text{ mod }m)$. Then $gx=(f(u),u\text{ mod }m)$. From the description of the graph $\Gamma(G,X_{m})$ one can see that $d(x,gx)\geq 2|u|\geq C_{1}\mathrm{diam}(X_{m})$. ∎
Our main motivation for considering the class $\mathcal{C}_{LD}(f)$ is the following:
@@
If $f:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is increasing and satisfies (1), then the class $\mathcal{C}_{LD}(f)$ is closed under finite graph products.
@@
Let $(G_{i})_{I}$ be a finite collection of groups in $\mathcal{C}_{LD}(f)$. For every $i\in I$ let $Y_{i}$ be a $G_{i}$-set such that the $G_{i}$-action on $Y_{i}$ has large displacement and growth at most $f(n)$. Let $S_{i}$ be a finite generating subset of $G_{i}$ and $C_{i}$ a constant as in Definition 2.9. Let $P=(G_{i})_{i\in I}^{c}$ be a graph product, generated by $S=\cup S_{i}$. Set $C=\max C_{i}$. For $i\in I$, let $\mathcal{F}_{i}$ be the set of $G_{i}$-orbits in $Y_{i}$, and $\mathcal{F}$ the disjoint union of the $\mathcal{F}_{i}$. We consider the subset of $\mathcal{G}(\mathcal{F},c)$, denoted by $Y$, consisting of the disjoint union of all $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$, where $q\geq 1$, $(X_{k})_{k\leq q}$ is a $c$-admissible $q$-tuple of elements of $\mathcal{F}$, and $e_{k},s_{k}$ are distinct elements of $X_{k}$ with the condition that $\mathrm{diam}(X_{k})\leq Cd(e_{k},s_{k})$ for every $k\leq q$. The group $P$ acts on $Y$, and every $P$-orbit in $Y$ is finite because all the members of $\mathcal{F}$ are finite. Moreover Lemma 2.8 ensures $\operatorname{vol}_{P,Y}(n)\preccurlyeq f(n)$. Hence to conclude the proof it suffices to see that the $P$-action on $Y$ has large displacement. Let $g$ be a non-trivial element of $P$. We repeat verbatim the argument in the proof of Lemma 2.6. We find a decomposition $g=\gamma_{q}\cdots\gamma_{1}$ and a $c$-admissible $(X_{k},e_{k},s_{k})_{k\leq q}$ such that $\gamma_{k}(e_{k})=s_{k}$ for every $k\leq q$ (so that $g(e_{1})=s_{q}$). The assumption that the $G_{i}$-action on $Y_{i}$ has large displacement allows to ensure that $\mathrm{diam}(X_{k})\leq Cd(e_{k},s_{k})$ for every $k\leq q$. Hence if we write $X=\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$, then $X\in Y$ and
@@ 2.4. The Grigorchuk group
In this section we prove that the Grigorchuk group belongs to $\mathcal{C}_{LD}(\mathrm{lin})$. We denote by $T=\{0,1\}^{*}$ the rooted tree of finite binary words, where each word $w$ is connected by an edge to $wx,x\in\{0,1\}$. Recall that the Grigorchuk group $G$ is the subgroup of the automorphism group $\operatorname{Aut}(T)$ generated by the set of automorphisms $S=\{a,b,c,d\}$ given by the recursive rules
$\begin{array}[]{lr}a(0w)=1w,&a(1w)=0w;\\ b(0w)=0a(w),&b(1w)=0c(w);\\ c(0w)=0a(w),&c(1w)=1d(w);\\ d(0w)=0w,&d(1w)=1b(w).\\ \end{array}$
We denote by $\Gamma_{n}$ the Schreier graph of the action of $G$ on the level $\{0,1\}^{n}$ of the tree. We recall basic properties of the structure of these graphs, see [BG00] for a more detailed description. Each $\Gamma_{n}$ is isometric to an interval of length $12^{n}-$ 1 in $\mathbb{Z}$ (with loops and multiple edges, which won’t be important for us). Along these segments, binary words are ordered following the Gray code ordering, namely each $w\in\{0,1\}$ is connected by an edge in $\Gamma_{n}$ to the word obtained by changing its first digit, and to the word obtained by changing the digit that follows the first appearance of 00 in $w$ (if the latter exists). Thus all words have exactly two neighbours, with the exception of the words $11\cdots 1$ 11 and $11\cdots 10$, which lie at the extreme points of $\Gamma_{n}$ (we shall picture this with $11\cdots 1$ 11 as the leftmost point and $11\cdots 10$ as the rightmost point). Let us denote by $d_{n}$ the associated distance on $\{0,1\}^{n}$. The map $p_{n}\colon\{0,1\}^{n}\to\{0,1\}^{n-1}$ that erases the last digit induces a covering map $p_{n}\colon\Gamma_{n}\to\Gamma_{n-1}$, which corresponds to folding $\Gamma_{n}$ around its middle edge.
@@
Let $G$ be the Grigorchuk group and retain the above notation. Then for every non-trivial element $g\in G$, there exists $n\geq 1$ and $w\in\{0,1\}^{n}$ such that $d_{n}(gw,w)\geq\frac{1}{8}\operatorname{diam}(\Gamma_{n})$. In particular $G$ belongs to $\mathcal{C}_{LD}(\mathrm{lin})$.
@@
Let $g\in G$ be non-trivial. Let $m$ be the smallest level containing active vertices. We can suppose that $G$ fixes $\{0,1\}^{n}$ for all $n\leq 3$, or the conclusion holds true trivially for $g$, since $\operatorname{diam}(\Gamma_{n})\leq 7$ for $n\leq 3$ and any point which is not fixed by $g$ is moved at distance at least 1. Hence $m\geq 3$. Choose $w\in\{0,1\}^{m}$ active, let $w_{0}$ be the projection of $w$ at level $m-3$, and $T_{0}$ be the finite rooted at $w_{0}$ and containing all its descendants up to 3 levels. Hence $w$ is a leaf of $T_{0}$, and all active vertices for $g$ on $T_{0}$ must be leaves, by minimality of $m$. It follows from Theorem 1 in [AS07] that at least another leaf of $T_{0}$, distinct from $w$, must be active. It follows that we can find two distinct active vertices $w,v\in\{0,1\}^{m}$ having the same projection at level $m-3$. Consider now the action on level $m+1$. Since $g$ fixes level $m$, it must preserves all fibers of the covering map $p_{m+1}\colon\Gamma_{m+1}\to\Gamma_{m}$, which consist each of two points, and it acts non-trivially on each of the two fibers $p_{m+1}^{-1}(w),p_{m+1}^{-1}(v)$. Let $I\subset\Gamma_{m}$ be the interval ending at the right-most points of length $2^{m-3}\geq\frac{1}{8}\operatorname{diam}(\Gamma_{m})$. For any vertex $u\notin I$, the two points in $p_{m+1}^{-1}(u)$ are at distance at least $2|I|\geq\frac{1}{8}\operatorname{diam}(\Gamma_{m+1})$ in $\Gamma_{m+1}$. On the other hand the projection map from $\Gamma_{m}\to\Gamma_{m-3}$ is injective on $I$, and since $w,v$ have the same image, they can’t both belong to $I$. Hence the desired conclusion follows, for $n=m+1$. ∎
2.5. The proof of Proposition 2
Note that Proposition 2 from the introduction follows from Proposition 2.13, together with the combination of Example 2.11, Proposition 2.12 and Proposition 2.14.

### v2

@@ 2.3. A subclass closed under graph products
In this section we prove that, under a mild assumption on the function $f$, a certain subclass of $\mathcal{C}(f)$ is closed under finite graph products (Proposition 2.13). The definition of this subclass is inspired by the work of the third author [Sal21], and Proposition 2.13 elaborates on [Sal21, Theorem 3]. In particular Proposition 2.13 provides an elementary proof of the result from [Sal21] that finitely generated right-angled Artin groups admit a faithful action with linear growth. (Actually, Theorem 3 in [Sal21] is stronger than that).
@@
Let $C_{1}$ be as in (1). Let $n\geq 1$. Let $x$ be a vertex in $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$, and let $\ell$ be such that $x$ belongs to $X_{\ell}$. Let $B(n)$ denote the ball of radius $n$ around $x$ in $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$. Let $X_{\ell-\ell_{1}},\ldots X_{\ell},\ldots X_{\ell+\ell_{2}}$ be the members of $(X_{k})_{k\leq q}$ that $B(n)$ intersects. Certainly we have
@@
Let $G$ be a finitely generated group, and $X$ a $G$-set. We say that the $G$-action on $X$ has orbits of controlled diameter if all the $G$-orbits in $X$ are finite, and if there is $C>0$ such that for every non-trivial element $g\in G$, there is a $G$-orbit $O$ in $X$ and $x\in O$ such that $gx\neq x$ and $\mathrm{diam}(\Gamma(G,O))\leq Cd(x,gx)$.
Here by $\Gamma(G,O)$ we mean the Schreier graph of the $G$-action on $O$ (which is a connected component in the larger graph $\Gamma(G,X)$), with respect to a finite symmetric generating set $S$. It is easily checked that this condition does not depend on the choice of $S$ (although the constant $C$ can depend on $S$). Note that the definition forces in particular the group $G$ to be residually finite.
@@
We denote by $\mathcal{C}_{\mathrm{CD}}(f)\subset\mathcal{C}(f)$ the class of groups admitting an action with orbits of controlled diameter and of growth at most $f(n)$.
@@
    Every faithful action of a finite group has orbits of controlled diameter.
@@
    The action of $\mathbb{Z}$ on $\bigsqcup_{n\geq 1}\mathbb{Z}/n\mathbb{Z}$ has orbits of controlled diameter. Hence $\mathbb{Z}$ belongs to $\mathcal{C}_{\mathrm{CD}}(\mathrm{lin})$.
@@
    If $G_{1},\ldots,G_{k}$ admit actions with orbits of controlled diameter respectively on $X_{1},\ldots,X_{k}$, then the action of $G_{1}\times\cdots\times G_{k}$ on $X_{1}\sqcup\cdots\sqcup X_{k}$ has orbits of controlled diameter. Hence each class $\mathcal{C}_{\mathrm{CD}}(f)$ is closed under direct products.
@@
For every $d\geq 1$ and $p\geq 2$, the wreath product $G=(\mathbb{Z}/p\mathbb{Z})\wr\mathbb{Z}^{d}$ belongs to $\mathcal{C}_{\mathrm{CD}}(f)$ for $f(n)=n^{d}$.
@@
We denote elements of $G$ as pairs $(f,u)$, with $f\colon\mathbb{Z}^{d}\to\mathbb{Z}/p\mathbb{Z}$ of finite support and $u\in\mathbb{Z}^{d}$. For $m\geq 2$, set $H_{m}=\{(f,u)\colon\quad\sum_{v\in(m\mathbb{Z})^{d}}f(v)=0,u\in(m\mathbb{Z})^{d}\}$. Then $H_{m}$ is a finite index subgroup of $G$, of index $pm^{d}$. Set $X_{m}:=G/H_{m}$ and $X=\sqcup_{m}X_{m}$. We check that the action of $G$ on $X$ has orbits of controlled diameter and satisfies $\operatorname{vol}_{G,X}(n)\simeq n^{d}$. Consider the standard generating set $S=\{(\delta_{0},0),(0,e_{i}),i=1,\dots d\}$, where $\delta_{0}\colon\mathbb{Z}^{d}\to\mathbb{Z}/p\mathbb{Z}$ takes the value 1 on 0 and 0 elsewhere, and $e_{1},\ldots,e_{d}$ is the standard basis of $\mathbb{Z}^{d}$. Note first that the map $G\to(\mathbb{Z}/p\mathbb{Z})\times(\mathbb{Z}/m\mathbb{Z})^{d}$ that sends $(f,u)\in G$ to $(\sum_{v\in m\mathbb{Z}^{d}}f(v),u\mod m)$ descends to a bijection of the coset space $G/H_{m}$ with $(\mathbb{Z}/p\mathbb{Z})\times(\mathbb{Z}/m\mathbb{Z})^{d}$, so that we may identify $X_{m}$ with $(\mathbb{Z}/p\mathbb{Z})\times(\mathbb{Z}/m\mathbb{Z})^{d}$. Under this identification, the action of $G$ on $X_{m}$ coincides with the standard wreath product action of the natural quotient $(\mathbb{Z}/p\mathbb{Z})\wr(\mathbb{Z}/m\mathbb{Z})^{d}$. Explicitly, the action of elements in the generating set $S$ is described as follows: the lamp generator $(\delta_{0},0)$ permutes cyclically $(\mathbb{Z}/p\mathbb{Z})\times\{0\}$ and acts trivially elsewhere; each element $(0,e_{i})$ maps $(r,u\text{ mod }m)$ to $(r,u+e_{i}\text{ mod }m)$. In particular the Schreier graph of the $G$-action on $X_{m}$ is isomorphic (ignoring loops) to the graph obtained taking a cycle of length $p$ and gluing to each point a copy of the standard Cayley graph of $(\mathbb{Z}/m\mathbb{Z})^{d}$. It follows that $\operatorname{vol}_{G,X}(n)\simeq n^{d}$, and that $C_{1}m\leq\mathrm{diam}(X_{m})\leq C_{2}m$ for some constants $C_{1},C_{2}$. To check that the action has orbits of controlled diameter, let $g=(f,v)$ be a non-trivial element of $G$. Suppose first that $v\neq 0$, and choose $m=2|v|$, where $|\cdot|$ denotes the standard word metric of $\mathbb{Z}^{d}$. Then, since the natural projection of $X_{m}$ to $(\mathbb{Z}/m\mathbb{Z})^{d}$ is equivariant, for every $x\in X_{m}$ we have $d(x,gx)\geq|v|\geq\frac{1}{2}C_{1}\mathrm{diam}(X_{m})$. Suppose now that $v=0$. Let $u$ be such that $f(u)\neq 0$ and with $|u|$ maximal, and choose again $m>2|u|$. Let $x\in X_{m}\simeq(\mathbb{Z}/p\mathbb{Z})\times(\mathbb{Z}/m\mathbb{Z})^{d}$ be the point $x=(0,u\text{ mod }m)$. Then $gx=(f(u),u\text{ mod }m)$. From the description of the graph $\Gamma(G,X_{m})$ one can see that $d(x,gx)\geq 2|u|\geq C_{1}\mathrm{diam}(X_{m})$. ∎
Our main motivation for considering the class $\mathcal{C}_{\mathrm{CD}}(f)$ is the following:
@@
If $f:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is increasing and satisfies (1), then the class $\mathcal{C}_{\mathrm{CD}}(f)$ is closed under finite graph products.
@@
Let $(G_{i})_{I}$ be a finite collection of groups in $\mathcal{C}_{\mathrm{CD}}(f)$. For every $i\in I$ let $Y_{i}$ be a $G_{i}$-set such that the $G_{i}$-action on $Y_{i}$ has orbits of controlled diameter and growth at most $f(n)$. Let $S_{i}$ be a finite generating subset of $G_{i}$ and $C_{i}$ a constant as in Definition 2.9. Let $P=\mathsf{GP}((G_{i})_{I};c)$ be a graph product, generated by $S=\bigcup S_{i}$. Set $C=\max_{i}C_{i}$.
For $i\in I$, let $\mathcal{F}_{i}$ be the set of $G_{i}$-orbits in $Y_{i}$, and $\mathcal{F}$ the disjoint union of the $\mathcal{F}_{i}$. We consider the subset of $\mathcal{G}(\mathcal{F},c)$, denoted by $Y$, consisting of the disjoint union of all $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$, where $q\geq 1$, $(X_{k})_{k\leq q}$ is a $c$-admissible $q$-tuple of elements of $\mathcal{F}$, and $e_{k},s_{k}$ are distinct elements of $X_{k}$ with the condition that $\mathrm{diam}(X_{k})\leq Cd(e_{k},s_{k})$ for every $k\leq q$. The group $P$ acts on $Y$, and every $P$-orbit in $Y$ is finite because all the members of $\mathcal{F}$ are finite. Moreover Lemma 2.8 ensures $\operatorname{vol}_{P,Y}(n)\preccurlyeq f(n)$. Hence to conclude the proof it suffices to see that the $P$-action on $Y$ has orbits of controlled diameter.
Let $g$ be a non-trivial element of $P$. We repeat verbatim the argument in the proof of Lemma 2.6. We find a decomposition $g=\gamma_{q}\cdots\gamma_{1}$ and a $c$-admissible $(X_{k},e_{k},s_{k})_{k\leq q}$ such that $\gamma_{k}(e_{k})=s_{k}$ for every $k\leq q$ (so that $g(e_{1})=s_{q}$). The assumption that the $G_{i}$-action on $Y_{i}$ has orbits of controlled diameter allows to ensure that $\mathrm{diam}(X_{k})\leq Cd(e_{k},s_{k})$ for every $k\leq q$. Hence if we write $X=\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$, then $X\in Y$ and
@@ 2.4. The Grigorchuk group
In this section we prove that the Grigorchuk group belongs to $\mathcal{C}_{\mathrm{CD}}(\mathrm{lin})$. We denote by $T=\{0,1\}^{*}$ the rooted tree of finite binary words, where each word $w$ is connected by an edge to $wx,x\in\{0,1\}$. Recall that the Grigorchuk group $G$ is the subgroup of the automorphism group $\operatorname{Aut}(T)$ generated by the set of automorphisms $S=\{a,b,c,d\}$ given by the recursive rules
[…]
We denote by $\Gamma_{n}$ the Schreier graph of the action of $G$ on the level $\{0,1\}^{n}$ of the tree. We recall basic properties of the structure of these graphs, see [BG00] for a more detailed description. Each $\Gamma_{n}$ is isometric to an interval of length $2^{n}-1$ in $\mathbb{Z}$ (with loops and multiple edges, which will not be important for us). Along these segments, binary words are ordered following the Gray code ordering, namely each $w\in\{0,1\}^{n}$ is connected by an edge in $\Gamma_{n}$ to the word obtained by changing its first digit, and to the word obtained by changing the digit that follows the first appearance of 0 in $w$ (if the latter exists). Thus all words have exactly two neighbours, with the exception of the words $11\cdots 11$ and $11\cdots 10$, which lie at the extreme points of $\Gamma_{n}$ (we shall picture this with $11\cdots 11$ as the leftmost point and $11\cdots 10$ as the rightmost point). Let us denote by $d_{n}$ the associated distance on $\{0,1\}^{n}$. The map $p_{n}\colon\{0,1\}^{n}\to\{0,1\}^{n-1}$ that erases the last digit induces a covering map $p_{n}\colon\Gamma_{n}\to\Gamma_{n-1}$, which corresponds to folding $\Gamma_{n}$ around its middle edge.
@@
Let $G$ be the Grigorchuk group and retain the above notation. Then for every non-trivial element $g\in G$, there exists $n\geq 1$ and $w\in\{0,1\}^{n}$ such that $d_{n}(gw,w)\geq\frac{1}{8}\operatorname{diam}(\Gamma_{n})$. In particular $G$ belongs to $\mathcal{C}_{\mathrm{CD}}(\mathrm{lin})$.
@@
Let $g\in G$ be non-trivial. Let $m$ be the smallest level containing active vertices. We can suppose that $G$ fixes $\{0,1\}^{n}$ for all $n\leq 3$, or the conclusion holds true trivially for $g$, since $\operatorname{diam}(\Gamma_{n})\leq 7$ for $n\leq 3$ and any point which is not fixed by $g$ is moved at distance at least 1. Hence $m\geq 3$. Choose $w\in\{0,1\}^{m}$ active, let $w_{0}$ be the projection of $w$ at level m3m-3 (i.e., the word obtained by removing the last three digits from $w$), and $T_{0}$ be the finite rooted at $w_{0}$ and containing all its descendants up to 3 levels. Hence $w$ is a leaf of $T_{0}$, and all active vertices for $g$ on $T_{0}$ must be leaves, by minimality of $m$. It follows from Theorem 1 in [AS07] that at least another leaf of $T_{0}$, distinct from $w$, must be active. It follows that we can find two distinct active vertices $w,v\in\{0,1\}^{m}$ having the same projection at level m3m-3. Consider now the action on level $m+1$. Since $g$ fixes level $m$, it must preserves all fibers of the covering map $p_{m+1}\colon\Gamma_{m+1}\to\Gamma_{m}$, which consist each of two points, and it acts non-trivially on each of the two fibers $p_{m+1}^{-1}(w),p_{m+1}^{-1}(v)$. Let $I\subset\Gamma_{m}$ be the interval ending at the right-most points of length $2^{m-3}\geq\frac{1}{8}\operatorname{diam}(\Gamma_{m})$. For any vertex $u\notin I$, the two points in $p_{m+1}^{-1}(u)$ are at distance at least $2|I|\geq\frac{1}{8}\operatorname{diam}(\Gamma_{m+1})$ in $\Gamma_{m+1}$. On the other hand the projection map from $\Gamma_{m}\to\Gamma_{m-3}$ is injective on $I$, and since $w,v$ have the same image, they cannot both belong to $I$. Hence the desired conclusion follows, for $n=m+1$. ∎
2.5. The proof of Proposition 1.3
Note that Proposition 1.3 from the introduction follows from Proposition 2.13, together with the fact that all groups in the list belong to $\mathcal{C}_{\mathrm{CD}}(\mathrm{lin})$, which has been established in Example 2.11, Proposition 2.12 and Proposition 2.14.